HETAV PATEL

Grade 11,

5, 7, 8, __, 9, 10, 11, 89, 12, 13, 14, 155. please help _ bitsat question

5, 7, 8, __, 9, 10, 11, 89, 12, 13, 14, 155.
 
please help _ bitsat question
 

Grade:12

2 Answers

ng29
9 years ago
answer is  33
 
5*8-7=33
11*9-10=89
12*14-13=155
kartheeek
9 years ago
ans is 5*9-7 33 bcos 11*9-10=89 which are 5,6,7,8 numbers in the sequence